OpenStudy (anonymous):

Identify the asymptotes of the graph of the function. Then Graph the function??? f(x)=1/x-5

OpenStudy (anonymous):

@dumbcow @ganeshie8 @SolomonZelman

OpenStudy (anonymous):

@Clueless_math @amistre64

OpenStudy (eric_d):

vertical asymtote=5

OpenStudy (anonymous):

@eric_d how did you get the answer?

OpenStudy (eric_d):

x-5=0

OpenStudy (anonymous):

@eric_d could you explain a little more? like step by step?

OpenStudy (eric_d):

http://www.purplemath.com/modules/asymtote.htm

OpenStudy (amistre64):

division by zero is undefined ...

OpenStudy (anonymous):

@amistre64 wait I don't get it?

OpenStudy (amistre64):

how does your material define an asymptote?

OpenStudy (anonymous):

@amistre64 a line is an asymptote of a graph if the graph gets closer to the line as x or y gets larger in absolute value.

OpenStudy (amistre64):

for oblique and horizontal asymptotes, that definition is not accurate since a graph can cross these types of asymptotes but rather even out to one side for larger values

OpenStudy (anonymous):

@amistre64 okay..

OpenStudy (amistre64):

oblique and horizontal asymps describe 'end' behaviour. whereas vertical asymps define midrange behaviours and usually end up blowing up towards infinity. \[\frac{1}{x-5}\] this has 1 vertical asymptote since /0, dividing by 0, has no value and that occurs when x-5=0 it has 1 horizontal asymptote when x is huge .... 1/HUGE is a very very small fraction. imagine if you had a cake that was sliced up into 1,000,000 pieces. and you got 1 slice. youd have practically no cake at all. As the number of slices increases ... the amount of cake you get tends to practically zero.

OpenStudy (amistre64):